solve the equation.
-9b + 1 + 7b = -3b + 6

Found 2 solutions by prabhjyot, zynan88:
Answer by prabhjyot(165)

-9b + 1 + 7b = -3b + 6
bringing like terms to one side
-9b + 7b + 3b = 6-1
1b= 5
b= 5

Answer by zynan88(4)

-9b + 1 + 7b = -3b + 6

-9b+7b+3b=6-1
1b=5
b=5/1
b=5